OpenStudy (anonymous):

Write equation of the line passing through those two points using the point-slope formula y-y1=m(x-x1) Show all your work remember to find the slope of the line first. The numbers are (4.2 and 4.7) and (7.2 and 7.9) Please help ME, i need an answer as fast as you can !

OpenStudy (anonymous):

You mean coordinates? Not numbers? Anyway. Points (4.2, 4.7) & (7.2, 7.9) M = y2-y1/ (x2-x1) = (7.9 - 4.7) / (7.2 - 4.2) = 3.2/3 M = 3.2/3 Now to find C, sub in one pair of points. e.g (4.2, 4.7) as 4.2 = x and 4.7 = y Equation of a straight line : y=mx+c 4.7=(3.2/3)(4.2) + c 4.7 = 4.48 + c c = 0.22 Now rewrite the equation of a straight line as y = (3.2/3)x + 0.22 I think that is right?

OpenStudy (anonymous):

yes, thank you verymuch man a big medal goes for you :)

OpenStudy (anonymous):

Okay and don't forget you can substitute in either coordinates, usually I just substitute in what ever doesn't contain a negative number. Just because it's easier. :)